Lori wants to give her daughter $25,000 in 8 years to start her own business. How much should Lori invest today, at an annual interest rate of 8%, compounded annually, to have $25,000 in 8 years? a. $12,802.95. b. $13,210.34. c. $13,347.70. d. $13,506.72.
Answer #1

PV = FV / (1 + r)n

= $25,000 / 1.088 = $25,000 / 1.8509 = $13,506.72

Hence, Option "D" is correct.